OpenStudy (anonymous):

Find the exact values of tan(7pi/6) and csc(7pi/6).

OpenStudy (s4sensitiveandshy):

unit circle ?

OpenStudy (s4sensitiveandshy):

reciprocals: \[\tan \theta =\frac{\sin \theta}{\cos \theta}\]\[\csc \theta= \frac{1}{\sin \theta}\]\[\sec \theta= \frac{1}{\cos \theta}\]

OpenStudy (s4sensitiveandshy):

what is sin at 7pi/6 angle ? and cos (7pi/6) ?

OpenStudy (dayakar):

7pi/6 = 7* 180/6 = 7*30 = 210 is it clear

OpenStudy (dayakar):

\[\tan \frac{7 \pi }{ 6 }= \tan 210\]

OpenStudy (dayakar):

tan 210 = tan (180+30) = tan 30 do u know tan 30 =?

OpenStudy (dayakar):

do u understand or not